Tags:
create new tag
view all tags

Shift 20101118

09:00 AA, LC, LD, LN

Plan

  • Same as before - load 10nov17_0005.dat set file (MW's optics tweaked by AA and LD), tune for low background and trajectory, do laser/ e beam timing, search for collisions.

Start

Damping ring changed from 10 bunches, 0.78 Hz to 1 bunch, 1.56Hz. Lasers on. Seed cleaned (power up from 280 to 320mW) and tuned to lock.

Okugi-san tuning extraction. Loss in extraction line high and it appears that one of the OTR screens was in the beam. After ~ 2hours tuning access taken to swap polarity of QM11FF and QM16FF, as required by MW's optics design (only QM16FF done on 16th Nov).

11.20am Magnets changed - load up 10nov17_0005.dat.

12.10pm Beam very unstable and hard to tune. Near resonance in the DR - possibly only good solution is complete retune of DR. AA, GW, MW tuning DR and extraction line.

2.15pm OTR0 screen still in beam for some reason. GW removed it manually, orbit changed and background dropped. GW, LD tuning extraction line. LD removed attenuation from signal (1V, now at 14dB, ~ 6000 counts). LC retimed PCs for regen - moves to longer build up times and slightly lower energies as the system warms up. Output ~ 14.5mJ.

2.15pm Linac laser up to 40% to give enough charge in extraction line. GW saved 10nov18_1425.dat, fallback position file. Background very sensitive to excursions at QF11 and QD12. GW tuning these down then correcting down the rest of the line.

4.20pm Access. LD, LN go to visually check the positions of the OTRs. Very difficult to tune for low background, leading to suspicion that they weren't actually out of the way. All OTRs sitting low (although OTR3 not bad) and have to be moved by Javier - remote control not working so he has to be in the accelerator.

4.45pm Beam back on. No beam making it to the damping ring.

5pm Access. Japanese want to check there is nothing wrong.

5.30pm Beam on. Tuning for background.

8.30pm Tuned to vertical beam sizes of ~ 30um at both MW1X and MW2X and horizontal beam sizes of ~ 400um at MW1X and ~ 150um at MW2X. Saved optics file 10nov18_2030.dat. Checked timing on normal scope, e beam alignment with notch in OTR screen.

9pm Collisions! Chamber 945 vertical 10 horizontal centre of peak, phase 2.75. After chamber alignment to e beam with notch we were right on top of the signal. Pointing of e beam very good with these optics which seems to have made the difference.

Wire Scanner Dimension Centre (mm) Width (Ám)
MW1X X 20.546▒0.118 449.2▒151.9
Y 53.209▒0.008 26.7▒9.0
MW2X X 34.172▒0.004 99.2▒4.5
Y 62.123▒0.004 30.6▒4.5

This gives an angular difference between this e beam and the OTR alignment laser used for a reference of Δx = 41 μrad and Δy = 38 μrad - much better in both directions at once than we have seen before.

Scanning vertically, horizontally and temporally to optimise the signal. Vertical and temporal very good, optimised horizontal position -150.

11.30pm LD tweaked QD18X from 31.214 to 31.464 A, improving signal maximum from ~ 4000 to 4500 counts.

12am LN trying to fit scans but python scripts had not been logging data - all files empty. GW archiving laserwire PVs so data can be retrieved and analysed.

1am Two scans with debugged scanning and logging data. Later fit shows a laserwire scan σ = 11.4 +/- 0.2 μm.

Edit | Attach | Watch | Print version | History: r6 < r5 < r4 < r3 < r2 | Backlinks | Raw View | Raw edit | More topic actions

Physics WebpagesRHUL WebpagesCampus Connect • Royal Holloway, University of London, Egham, Surrey TW20 0EX; Tel/Fax +44 (0)1784 434455/437520

Topic revision: r6 - 21 Nov 2010 - LauraCorner

 
This site is powered by the TWiki collaboration platform Powered by PerlCopyright © 2008-2021 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding RHUL Physics Department TWiki? Send feedback